# Improved Generalization Bounds for Communication Efficient Federated Learning

### v1

@@ 1 Introduction
The primary purpose of communication in federated learning is to periodically aggregate local models to reduce the consensus distance among clients. This practice helps maintain the overall optimization process on a trajectory toward global optimization. It is important to note that when the consensus distance among clients becomes substantial, the convergence rate reduces. This occurs as individual clients gradually veer towards their respective local optima without being synchronized with models from other clients. This issue is amplified when the data distribution among clients is non-iid. It has been demonstrated that the consensus distance is correlated to (i) the randomness in each client’s own dataset, which causes variation in consecutive local gradients, as well as (ii) the dissimilarity in loss functions among clients due to non-iidness (Stich and Karimireddy,, 2019; Gholami and Seferoglu,, 2024). More specifically, the consensus distance at iteration $t$ is defined as $\frac{1}{K}\sum_{k=1}^{K}\|\hat{{\boldsymbol{\theta}_{t}}}-\boldsymbol{\theta}_{k,t}\|^{2}$, where $\hat{{\boldsymbol{\theta}_{t}}}=\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{\theta}_{k,t}$, $K$ is the number of clients, $\boldsymbol{\theta}_{k,t}$ is the local model at client $k$ at iteration $t$, and .$\|.\|^{2}$ is squared $l_{2}$ norm. Note that the consensus distance goes to zero when global aggregation is performed at each communication round. This makes the communication of models between clients and the parameter server crucial, but this introduces significant communication overhead. Our goal in this paper is to reduce the communication overhead of federated learning through the following contributions.
Contribution I: Improved Generalization Error Bound. The generalization error of a learning model is defined as the difference between the model’s empirical and population risks. (We provide a mathematical definition in Section 3). Existing approaches for training models mostly minimize the empirical risk or its variants. However, a small population risk is desired showing how well the model performs in the test phase as it denotes the loss that occurs when new samples are randomly drawn from the distribution. Note that a small empirical risk and a reduced generalization error correspond to a low population risk. Thus, there is an increasing interest in establishing an upper limit for the generalization error and understanding the underlying factors that affect the generalization error. The generalization error analysis is also important to quantitatively assess the generalization characteristics of trained models, provide reliable guarantees concerning their anticipated performance quality, and design new models and systems.
@@
Contribution II: Representation Learning Interpretation. Recent studies have demonstrated that the concept of representation learning is a promising approach to reduce the communication cost of federated learning (Collins et al.,, 2021). This is achieved by leveraging the shared representations that exist in the datasets of all clients. For example, let us consider a federated learning application for image classification, where different clients have datasets of different animals. Despite each client having a different dataset (one client has dog images, another has cat images, etc.), these images usually have common features such as an eye/ear shape. These shared features, typically extracted in the same way for different types of animals, require consistent layers of a neural network to extract them, whether the animal is a dog or a cat. As a result, these layers demonstrate similarity (i.e., less variation) across clients even when the datasets are non-iid. This implies that the consensus distance for this part of the model (feature extraction) is likely smaller. Based on these observations, our key idea is to reduce the aggregation frequency of the layers that show high similarity, where these layers are updated locally between consecutive aggregations. This approach would reduce the communication cost of federated learning as some layers are aggregated, hence their parameters are exchanged, less frequently. This makes it crucial to determine the layers that show high similarity. The next example scratches the surface of the problem for a toy example.
@@
We consider a federated learning setup of five clients with a central parameter server to train a ResNet-$20$ (He et al.,, 2015) on a heterogeneous partition of CIFAR-$10$ dataset (Krizhevsky,, 2009). We use Federated Averaging (FedAvg) (McMahan et al.,, 2016) as an aggregation algorithm as it stands as the dominant algorithm in federated learning. We applied FedAvg with $50$ local steps prior to each averaging step, denoted as $\tau=50$. Non-iidness is introduced by allocating $2$ classes to each client. Finally, we evaluate the quantity of the average consensus distance for each model layer during the optimization in Fig. 1. It is clear that the initial layers have smaller consensus distance as compared to the final layers. This is due to initial layers’ role in extracting representations from input data and their higher similarity across clients.
@@
The above example indicates that initial layers show higher similarity, so they can be aggregated less frequently. Additionally, several empirical studies (Reddi et al.,, 2021; Yu et al.,, 2020) show that federated learning with multiple local updates per round learns a generalizable representation and is unexpectedly successful in non-iid federated learning. These studies encourage us to delve deeper into investigating how local updates and model aggregation frequency affect the model’s representation extractor in terms of its generalization.
@@ 2 Related Work
There has been an increasing interest in distributed learning recently, largely driven by Federated Learning, which focuses on the analysis of decentralized algorithms utilizing Local SGD. Several studies have highlighted that these algorithms achieve convergence to a global optimum or a stationary point of the overall objective, particularly in convex or non-convex scenarios (Stich and Karimireddy,, 2019; Stich,, 2018; Gholami and Seferoglu,, 2024; Lian et al.,, 2018; Kairouz et al.,, 2019). However, it is widely accepted that communication cost is the major bottleneck for this technique in large-scale optimization applications (Konecný et al.,, 2016; Lin et al.,, 2017). To tackle this issue, two primary strategies are put forth: the utilization of mini-batch parallel SGD, and the adoption of Local SGD. These approaches aim to enhance the equilibrium between computation and communication. Woodworth et al., 2020b ; Woodworth et al., 2020a attempt to theoretically capture the distinction to comprehend under what circumstances Local SGD outperforms minibatch SGD.
@@ 3.1 Preliminaries and Notation
We consider that we have $K$ clients/nodes in our system, and each node has its own portion of the dataset. For example, node $k$ has a local dataset $\boldsymbol{S}_{k}=\{\boldsymbol{z}_{{k},{1}},...,\boldsymbol{z}_{{k},{n_{k}}}\}$, where $\boldsymbol{z}_{{k},{i}}=(\boldsymbol{x}_{{k},{i}},\boldsymbol{y}_{{k},{i}})$ is drawn from a distribution $\mathcal{D}_{k}$ over $\mathcal{X}\times\mathcal{Y}$ , where $\mathcal{X}$ is the input space and $\mathcal{Y}$ is the label space. The size of the local dataset at node $k$ is $n_{k}$. The dataset across all nodes is defined as $\boldsymbol{S}=\{\boldsymbol{S}_{1},...,\boldsymbol{S}_{K}\}$. Data distribution across the nodes could be independent and identically distributed (iid) or non-iid. In iid setting, we assume that $\mathcal{D}_{1}=...=\mathcal{D}_{K}=\mathcal{D}$ holds. On the other hand, non-iid setting covers all possible distributions and cases, where $\mathcal{D}_{1}=...=\mathcal{D}_{K}=\mathcal{D}$ does not hold.
@@ 3.2 Generalization Error
The expected generalization is expressed as $\operatorname{\mathbb{E}}_{\boldsymbol{S}}\Delta_{\mathcal{A}}(\boldsymbol{S})$, where $\operatorname{\mathbb{E}}_{\boldsymbol{S}}[.]=\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}[.]$ is used for the sake of notation convenience.
@@ 3.3 Federated Learning
The local models are usually trained using stochastic gradient descent (SGD) at each node. To reduce the communication cost needed between the nodes and the parameter server, each node executes multiple SGD steps using its local data after receiving an aggregated model from the parameter server. To be precise, we define the aggregated model parameters at round $r$ as $\hat{{\boldsymbol{\theta}}}_{r}$. Specifically, upon receiving $\hat{{\boldsymbol{\theta}}}_{r}$, node $k$ computes
@@
for $t=0,\dots,\tau-1$, where $\tau$ is the number of local SGD steps, $\boldsymbol{\theta}_{k,r,0}$ is defined as $\boldsymbol{\theta}_{k,r,0}=\hat{{\boldsymbol{\theta}}}_{r}$, $\eta$ is the learning rate, $\mathcal{B}_{k,r,t}$ is the batch of samples used in local step $t$ of round $r$ in node $k$, $\nabla$ is the gradient, and |.||.|| . | shows the size of a set. Upon completing the local steps in round $r$, each node transmits $\boldsymbol{\theta}_{k,r}=\boldsymbol{\theta}_{k,r,\tau}$ to the parameter server to calculate $\hat{{\boldsymbol{\theta}}}_{r+1}$ as in (4).
@@ 4.1 One-Round Generalization Bound
In the following lemma, we determine the generalization bound for one round of FedAvg.
@@
Let $l(M_{\boldsymbol{\theta}},\boldsymbol{z})$ be $\mu$-strongly convex and $L$-smooth in $M_{\boldsymbol{\theta}}$, $M_{{\boldsymbol{\theta}}_{k}}=\mathcal{A}_{k}(\boldsymbol{S}_{k})$ represents the model obtained from Empirical Risk Minimization (ERM) algorithm on local dataset $\boldsymbol{S}_{k}$, i.e., $M_{{\boldsymbol{\theta}}_{k}}=\arg\min_{M}\sum_{i=1}^{n_{k}}l(M,\boldsymbol{z}_{{k},{i}})$, and $M_{\hat{{\boldsymbol{\theta}}}}=\mathcal{A}(\boldsymbol{S})$ is the model after one round of FedAvg ($\hat{{\boldsymbol{\theta}}}=\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\boldsymbol{\theta}_{k}$). Then, the expected generalization error is
$\operatorname{\mathbb{E}}$ ${}_{\boldsymbol{S}}\Delta_{\mathcal{A}}(\boldsymbol{S})\leq\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\bigg{[}\frac{L\mathcal{K}(k)^{2}}{\mu}\underbrace{\operatorname{\mathbb{E}}_{\boldsymbol{S}_{k}}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})}_{\text{Expected local generalization}}+2\sqrt{\frac{L}{\mu}}\mathcal{K}(k)\bigg{(}\underbrace{\operatorname{\mathbb{E}}_{\boldsymbol{S}}\delta_{k,\mathcal{A}}(\boldsymbol{S})}_{\text{Expected non-iidness}}\underbrace{\operatorname{\mathbb{E}}_{\boldsymbol{S}_{k}}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})}_{\text{Expected local generalization}}\bigg{)}^{\frac{1}{2}}\bigg{]},$
@@
The generalization error results in these cases are affected by substituting $\frac{1}{\hat{K}}$ and $\frac{\mathcal{K}(k)K}{\hat{K}}$ instead of $\mathcal{K}(k)$ in (4.1) for cases I and II, respectively. The detailed proof is provided in Appendix C.
Discussion. Note that there are two terms in the generalization error bound: (i) local generalization of each client that shows more generalizable local models lead to a better generalization of the aggregated model, (ii) non-iidness of each client which deteriorates generalization. Theorem 4.1 reveals a factor of $\mathcal{K}(k)^{2}$ for the first term, which is the sole term in the iid setting. For example, in the uniform case ($\mathcal{K}(k)=\frac{1}{K}$), we will observe an improvement with a factor of $\frac{1}{K^{2}}$ for the iid case. This represents an enhancement compared to the state of the art Barnes et al., 2022b , which only demonstrates a factor of $\frac{1}{K}$. As a result, after each averaging process carried out by the central parameter server, the generalization error is reduced by a factor of $\mathcal{K}(k)^{2}$ in iid case.
@@ 4.2 $R-$Round Generalization Bound
Now we turn our attention to a more complicated scenario; $R$-round FedAvg. A similar $R$-round generalization bound analysis is considered in Barnes et al., 2022b , but without representation learning, which is crucial in our proposed federated learning mechanism in Section 6.
@@
Note that the expression in (7) differs slightly from the end-to-end generalization error that would be obtained by considering the final model $M_{\hat{\boldsymbol{\theta}}_{R}}$ and the entire dataset $\boldsymbol{S}$. More specifically, (7) is an average of the generalization errors measured at each round. We anticipate that the generalization error diminishes with the increasing number of data samples, so this generalization error definition yields to a more cautious upper limit and serves as a sensible measure. The next theorem characterizes the expected generalization error bounds for $R-$Round FedAvg in iid and non-iid settings.
@@
Let $l(M_{\boldsymbol{\theta}},\boldsymbol{z})$ be $\mu$-strongly convex and $L$-smooth in $M_{\boldsymbol{\theta}}$. Local models at round $r$ are calculated by doing $\tau$ local gradient descent steps and the local gradient variance is bounded by $\sigma^{2}$ ($\frac{1}{|Z_{k,r}|}\sum_{i\in Z_{k,r}}\|\nabla l(M_{\boldsymbol{\theta}},\boldsymbol{z}_{{k},{i}})-\frac{1}{|Z_{k,r}|}\sum_{i\in Z_{k,r}}\nabla l(M_{\boldsymbol{\theta}},\boldsymbol{z}_{{k},{i}})\|^{2}\leq\sigma^{2}$). The aggregated model at round $r$ is $M_{\hat{{\boldsymbol{\theta}}}_{r}}$ is obtained by performing FedAvg and where the data points used in round $r$ (i.e., $Z_{k,r}$) are sampled without replacement. Then the average generalization error bound is
$\operatorname{\mathbb{E}}_{\boldsymbol{S}}\Delta_{FedAvg}(\boldsymbol{S})\leq\frac{1}{R}\sum_{r=1}^{R}$ $\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\bigg{[}\frac{2L\mathcal{K}(k)^{2}}{\mu}A+\sqrt{\frac{8L}{\mu}}\mathcal{K}(k)(AB)^{\frac{1}{2}}\bigg{]}$ (8)

### v3

@@ 1 Introduction
The primary purpose of communication in federated learning is to periodically aggregate local models to reduce the consensus distance among clients. This practice helps maintain the overall optimization process on a trajectory toward global optimization. It is important to note that when the consensus distance among clients becomes substantial, the convergence rate reduces. This occurs as individual clients gradually veer towards their respective local optima without being synchronized with the models from other clients. This issue is amplified when the data distribution among clients is non-iid. It has been demonstrated that the consensus distance is correlated to (i) the randomness in each client’s own dataset, which causes variation in consecutive local gradients, as well as (ii) the dissimilarity in loss functions among clients due to non-iidness (Stich and Karimireddy,, 2019; Gholami and Seferoglu,, 2024). More specifically, the consensus distance at iteration $t$ is defined as $\frac{1}{K}\sum_{k=1}^{K}\|\hat{{\boldsymbol{\theta}_{t}}}-\boldsymbol{\theta}_{k,t}\|^{2}$, where $\hat{{\boldsymbol{\theta}_{t}}}=\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{\theta}_{k,t}$, $K$ is the number of clients, $\boldsymbol{\theta}_{k,t}$ is the local model at client $k$ at iteration $t$, and $\|\cdot\|^{2}$ is squared $l_{2}$ norm. Note that the consensus distance goes to zero when global aggregation is performed at each communication round. This makes the communication of models between clients and the parameter server crucial, but this introduces significant communication overhead. Our goal in this paper is to reduce the communication overhead of federated learning through the following contributions.
Contribution I: Improved Generalization Error Bound. The generalization error of a learning model is defined as the difference between the model’s empirical risk and population risks. (We provide a mathematical definition in Section 3). Existing approaches for training models mostly minimize the empirical risk or its variants. However, a small population risk is desired showing how well the model performs in the test phase as it denotes the loss that occurs when new samples are randomly drawn from the distribution. Note that a small empirical risk and a reduced generalization error correspond to a low population risk. Thus, there is an increasing interest in establishing an upper limit for the generalization error and understanding the underlying factors that affect the generalization error. The generalization error analysis is also important to quantitatively assess the generalization characteristics of trained models, provide reliable guarantees concerning their anticipated performance quality, and design new models and systems.
@@
Contribution II: Representation Learning Interpretation. Recent studies have demonstrated that the concept of representation learning is a promising approach to reducing the communication cost of federated learning (Collins et al.,, 2021). This is achieved by leveraging the shared representations that exist in the datasets of all clients. For example, let us consider a federated learning application for image classification, where different clients have datasets of different animals. Despite each client having a different dataset (one client has dog images, another has cat images, etc.), these images usually have common features such as an eye/ear shape. These shared features, typically extracted in the same way for different types of animals, require consistent layers of a neural network to extract them, whether the animal is a dog or a cat. As a result, these layers demonstrate similarity (i.e., less variation) across clients even when the datasets are non-iid. This implies that the consensus distance for this part of the model (feature extraction) is likely smaller. Based on these observations, our key idea is to reduce the aggregation frequency of the layers that show high similarity, where these layers are updated locally between consecutive aggregations. This approach would reduce the communication cost of federated learning as some layers are aggregated, hence their parameters are exchanged, less frequently. The next example scratches the surface of the problem for a toy example.
@@
We consider a federated learning setup of five clients with a central parameter server to train a ResNet-$20$ (He et al.,, 2015) on a heterogeneous partition of CIFAR-$10$ dataset (Krizhevsky,, 2009). We use Federated Averaging (FedAvg) (McMahan et al.,, 2016) as an aggregation algorithm since it is the dominant algorithm in federated learning. We applied FedAvg with $50$ local steps prior to each averaging step, denoted as $\tau=50$. Non-iidness is introduced by allocating $2$ classes to each client. Finally, we evaluate the quantity of the average consensus distance for each model layer during the optimization in Fig. 1. It is clear that the initial layers have smaller consensus distance as compared to the final layers. This is due to initial layers’ role in extracting representations from input data and their higher similarity across clients.
@@
The above example indicates that initial layers show higher similarity, so they can be aggregated less frequently. Additionally, several empirical studies (Reddi et al.,, 2021; Yu et al.,, 2020) show that federated learning with multiple local updates per round learns a generalizable representation and is unexpectedly successful in non-iid settings. These studies encourage us to delve deeper into investigating how local updates and model aggregation frequency affect the model’s representation extractor in terms of its generalization.
@@ 2 Related Work
There has been increasing interest in distributed learning recently, largely driven by Federated Learning. Several studies have highlighted that these algorithms achieve convergence to a global optimum or a stationary point of the overall objective, particularly in convex or non-convex scenarios (Stich and Karimireddy,, 2019; Stich,, 2018; Gholami and Seferoglu,, 2024; Lian et al.,, 2018; Kairouz et al.,, 2019). However, it is widely accepted that communication cost is the major bottleneck for these techniques in large-scale optimization applications (Konecný et al.,, 2016; Lin et al.,, 2017). To tackle this issue, two primary strategies are put forth: the utilization of mini-batch parallel SGD, and the adoption of Local SGD. These approaches aim to enhance the equilibrium between computation and communication. Woodworth et al., 2020b ; Woodworth et al., 2020a attempt to theoretically capture the distinction to comprehend under what circumstances Local SGD outperforms minibatch SGD.
@@ 3.1 Preliminaries and Notation
We consider that we have $K$ clients/nodes in our system, and each node has its own portion of the dataset. For example, node $k$ has a local dataset $\boldsymbol{S}_{k}=\{\boldsymbol{z}_{{k},{1}},...,\boldsymbol{z}_{{k},{n_{k}}}\}$, where $\boldsymbol{z}_{{k},{i}}=(\boldsymbol{x}_{{k},{i}},\boldsymbol{y}_{{k},{i}})$ is drawn from a distribution $\mathcal{D}_{k}$ over $\mathcal{X}\times\mathcal{Y}$ , where $\mathcal{X}$ is the input space and $\mathcal{Y}$ is the label space. We consider $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}\subseteq\mathbb{R}$ The size of the local dataset at node $k$ is $n_{k}$. The dataset across all nodes is defined as $\boldsymbol{S}=\{\boldsymbol{S}_{1},...,\boldsymbol{S}_{K}\}$. Data distribution across the nodes could be independent and identically distributed (iid) or non-iid. In iid setting, we assume that $\mathcal{D}_{1}=...=\mathcal{D}_{K}=\mathcal{D}$ holds. On the other hand, non-iid setting covers all possible distributions and cases, where $\mathcal{D}_{1}=...=\mathcal{D}_{K}=\mathcal{D}$ does not hold.
@@ 3.2 Generalization Error
The expected generalization error is expressed as $\operatorname{\mathbb{E}}_{\boldsymbol{S}}\Delta_{\mathcal{A}}(\boldsymbol{S})$, where $\operatorname{\mathbb{E}}_{\boldsymbol{S}}[\cdot]=\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}[\cdot]$ is used for the sake of notation convenience.
@@ 3.3 Federated Learning
The local models are usually trained using stochastic gradient descent (SGD) at each node. To reduce the communication cost needed between the nodes and the parameter server, each node executes multiple SGD steps using its local data after receiving an aggregated model from the parameter server. To be precise, we have the aggregated model parameters at round $r$ as $\hat{{\boldsymbol{\theta}}}_{r}$. Specifically, upon receiving $\hat{{\boldsymbol{\theta}}}_{r}$, node $k$ computes
@@
for $t=0,\dots,\tau-1$, where $\tau$ is the number of local SGD steps, $\boldsymbol{\theta}_{k,r,0}$ is defined as $\boldsymbol{\theta}_{k,r,0}=\hat{{\boldsymbol{\theta}}}_{r}$, $\eta$ is the learning rate, $\mathcal{B}_{k,r,t}$ is the batch of samples used in local step $t$ of round $r$ in node $k$, $\nabla$ is the gradient, and ||$\cdot$|| ⋅ | shows the size of a set. Upon completing the local steps in round $r$, each node transmits $\boldsymbol{\theta}_{k,r}=\boldsymbol{\theta}_{k,r,\tau}$ to the parameter server to calculate $\hat{{\boldsymbol{\theta}}}_{r+1}$ as in (4).
@@ 4.1 One-Round Generalization Bound
In the following theorem, we determine the generalization bound for one round of FedAvg.
@@
Let $l(M_{\boldsymbol{\theta}},\boldsymbol{z})$ be $\mu$-strongly convex and $L$-smooth in $M_{\boldsymbol{\theta}}$. $M_{{\boldsymbol{\theta}}_{k}}=\mathcal{A}_{k}(\boldsymbol{S}_{k})$ represents the model obtained from Empirical Risk Minimization (ERM) algorithm on local dataset $\boldsymbol{S}_{k}$, i.e., $M_{{\boldsymbol{\theta}}_{k}}=\arg\min_{M}\sum_{i=1}^{n_{k}}l(M,\boldsymbol{z}_{{k},{i}})$, and $M_{\hat{{\boldsymbol{\theta}}}}=\mathcal{A}(\boldsymbol{S})$ is the model after one round of FedAvg ($\hat{{\boldsymbol{\theta}}}=\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\boldsymbol{\theta}_{k}$). Then, the expected generalization error is
$\operatorname{\mathbb{E}}$ ${}_{\boldsymbol{S}}\Delta_{\mathcal{A}}(\boldsymbol{S})\leq\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\bigg{[}\frac{L\mathcal{K}(k)^{2}}{\mu}\underbrace{\operatorname{\mathbb{E}}_{\boldsymbol{S}_{k}}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})}_{\text{Expected local generalization}}+2\sqrt{\frac{L}{\mu}}\mathcal{K}(k)\bigg{(}\underbrace{\operatorname{\mathbb{E}}_{\boldsymbol{S}}\delta_{k,\mathcal{A}}(\boldsymbol{S})}_{\text{Expected non-iidness}}\underbrace{\operatorname{\mathbb{E}}_{\boldsymbol{S}_{k}}\Delta_{\mathcal{A}_{k}}(\boldsymbol{S}_{k})}_{\text{Expected local generalization}}\bigg{)}^{\frac{1}{2}}\bigg{]},$ (6)
@@
The generalization error results in these cases are affected by substituting $\frac{1}{\hat{K}}$ and $\frac{\mathcal{K}(k)K}{\hat{K}}$ instead of $\mathcal{K}(k)$ in (6) for cases I and II, respectively. The detailed proof is provided in Appendix C.
Discussion. Note that there are two terms in the generalization error bound: (i) local generalization of each client that shows more generalizable local models lead to a better generalization of the aggregated model, (ii) non-iidness of each client which deteriorates generalization. Theorem 4.1 reveals a factor of $\mathcal{K}(k)^{2}$ for the first term, which is the sole term in the iid setting. For example, in the uniform case ($\mathcal{K}(k)=\frac{1}{K}$), we will observe an improvement with a factor of $\frac{1}{K^{2}}$ for the iid case. This represents an enhancement compared to the state of the art (Barnes et al., 2022b, ; Sun et al.,, 2023; Yagli et al.,, 2020), which only demonstrates a factor of $\frac{1}{K}$. As a result, after the averaging process carried out by the central parameter server, the generalization error is reduced by a factor of $\mathcal{K}(k)^{2}$ in iid case.
@@ 4.2 $R-$Round Generalization Bound
Note that the expression in (8) differs slightly from the end-to-end generalization error that would be obtained by considering the final model $M_{\hat{\boldsymbol{\theta}}_{R}}$ and the entire dataset $\boldsymbol{S}$. More specifically, (8) is an average of the generalization errors measured at each round, similar to Barnes et al., 2022b ). We anticipate that the generalization error diminishes with the increasing number of data samples, so this generalization error definition yields to a more cautious upper limit and serves as a sensible measure. The next theorem characterizes the expected generalization error bounds for $R-$Round FedAvg in iid and non-iid settings.
@@
Let $l(M_{\boldsymbol{\theta}},\boldsymbol{z})$ be $\mu$-strongly convex and $L$-smooth in $M_{\boldsymbol{\theta}}$. Local models at round $r$ are calculated by doing $\tau$ local gradient descent steps (5), and the local gradient variance is bounded by $\sigma^{2}$, i.e., $\operatorname{\mathbb{E}}_{\boldsymbol{z}\sim\mathcal{D}_{k}}\|\nabla l(M_{\boldsymbol{\theta}},\boldsymbol{z})-\operatorname{\mathbb{E}}_{\boldsymbol{z}\sim\mathcal{D}_{k}}\nabla l(M_{\boldsymbol{\theta}},\boldsymbol{z})\|^{2}\leq\sigma^{2}$. The aggregated model at round $r$, $M_{\hat{{\boldsymbol{\theta}}}_{r}}$, is obtained by performing FedAvg, and the data points used in round $r$ (i.e., $Z_{k,r}$) are sampled without replacement. Then the average generalization error bound is
$\operatorname{\mathbb{E}}_{\boldsymbol{S}}\Delta_{FedAvg}(\boldsymbol{S})\leq\frac{1}{R}\sum_{r=1}^{R}$ $\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\bigg{[}\frac{2L\mathcal{K}(k)^{2}}{\mu}A+\sqrt{\frac{8L}{\mu}}\mathcal{K}(k)(AB)^{\frac{1}{2}}\bigg{]},$ (9)
@@
Proof: The proof of Theorem 4.3 is provided in Appendix B of the supplementary materials. $\Box$